1. If you would like to accumulate RM 8,000 over the next 5 years, how much you must deposit at the end of each six months, starting six months from now, given a 6 percent interest rate and semi-annual compounding?



2. What is the effective annual percentage rate (EAR) of 12 percent compounded monthly?



3. Why do you think that the application of time value of money is important for financial management? Discuss (not more than 500 words)

1
Expert's answer
2021-12-27T08:58:45-0500

Answer 1:

Future value (FV) = RM 8,000

Time period = 5 years

Frequency of compounding = 2 per year

n = 10

Interest rate = 6% p.a.

Semi-annual compounding (i) = 3%


"PMT=\frac{FV\times i}{[(1+i)^n -1]}"


"PMT=\frac{8000\times 0.03}{[(1+0.03)^10 -1]}=697.84"


This means there is need to deposit RM 697.84 at the end of each six month to accumulate RM 8,000 over the next 5 years.


Answer 2:

The effective annual percentage rate (EAR) of 12 percent compounded monthly is:

"EAR=(1+\frac{i}{n})^n-1"


"EAR=(1+\frac{0.12}{12})^12-1"


"EAR=1.1268-1"


"EAR=0.1268"


This mean EAR is 12.68%.


Answer 3:

There is greater application of time value of money concept for financial management. It is becauase this concept assists financial managers and investors to make a more informed decision. In financial management, time value of money has application in many areas like stock valuation, capital budgeting etc. This concept allows financial managers to evaluate potential investment projects by determining present value of their estimated future cash flows. Based on present value, managers can find out which capital investment project is most economical for the company.

Time value of money also has application in considering inflation, risk and investment opportunity before making investment decisions. Most of the financial management decisions involve costs and benefits. The concept of time value of money allows to perform cost and benefit analysis more effectively to make best suitable investment decisions.